### v4

# DCNv3: Towards Next Generation Deep Cross Network for Click-Through Rate Prediction

Abstract.
Deep & Cross Network and its derivative models have become an important paradigm in click-through rate (CTR) prediction due to their effective balance between computational cost and performance. However, these models face four major limitations: (1) while most models claim to capture high-order feature interactions, they often do so implicitly and non-interpretably through deep neural networks (DNN), which limits the trustworthiness of the model’s predictions; (2) the performance of existing explicit feature interaction methods is often weaker than that of implicit DNN, undermining their necessity; (3) many models fail to adaptively filter noise while enhancing the order of feature interactions; (4) the fusion methods of most models cannot provide suitable supervision signals for their different interaction methods.
To address the identified limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and Shallow & Deep Cross Network (SDCNv3). These models ensure interpretability in feature interaction modeling while exponentially increasing the order of feature interactions to achieve genuine Deep Crossing rather than just Deep & Cross. Additionally, we employ a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we use a simple yet effective loss weight calculation method called Tri-BCE to provide appropriate supervision signals. Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. The code, running logs, and detailed hyperparameter configurations are available at: https://anonymous.4open.science/r/DCNv3-E352.
Feature Interaction, Cross Network, Recommender Systems, CTR Prediction
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Figure 1. Comparison among DCNv3, SDCNv3, and other models in terms of network parameter number, AUC, and running time on Criteo dataset. $\star$ represents models using only explicit feature interactions, and $\circ$ represents models integrating implicit feature interactions. The graphic area represents the running time per epoch for each model (a larger area indicates a longer time, and vice versa).
Click-through rate (CTR) prediction is an essential part of industrial recommender systems (Zhu et al., 2021, 2022a). It uses user profiles, item attributes, and context to predict the probability of user-item interactions, thereby providing a better user experience and increasing the profitability of the recommender system (Wang et al., 2021b; Li et al., 2023; Chen et al., 2021; Guo et al., 2017). Most feature interaction-based CTR prediction models follow the paradigm proposed by DCN (Wang et al., 2017, 2021b), which aims to construct both explicit and implicit feature interactions and fuse the predictions of different interaction information to enhance interpretability and accuracy.
Despite the effectiveness of the current CTR paradigm, there are limitations to overcome:
  • Lack of interpretability. As shown in Figure 1, most models integrate deep neural networks (Covington et al., 2016) (DNN) to model implicit high-order feature interactions and achieve AUC performance between 81.3 and 81.5. This demonstrates the effectiveness of implicit feature interactions. However, implicit feature interactions lack interpretability (Marcus, 2018), which significantly reduces the trustworthiness of deep CTR model predictions.
  • Limited necessity for explicit interactions. As observed in Figure 1, most models (e.g., CIN (Lian et al., 2018)) using only explicit feature interactions achieve AUC performance below 81.3, while DNN achieves a better AUC of 81.4. This indicates that the performance of most explicit modeling methods is weaker than that of implicit DNN, which undoubtedly undermines the necessity of integrating explicit and implicit feature interactions. Therefore, FinalMLP (Mao et al., 2023) attempts to model features implicitly in a dual manner, discarding traditional explicit interaction methods, and thereby achieving state-of-the-art performance.
  • Ineffective noise filtering capability. Many studies (Liu et al., 2020; Zhu et al., 2023a; Deng et al., 2021; Wang et al., 2023a) point out that CTR models contain a significant amount of redundant feature interactions and noise, especially in higher-order feature interactions. Consequently, most CTR models (Song et al., 2019; Wang et al., 2021b; Lian et al., 2018) are built with only two or three network layers, abandoning the explicit capture of effective higher-order feature interaction information. Meanwhile, filtering noise for the model often incurs additional computational costs, which can lead to longer training and inference times, potentially offsetting the benefits gained from improved model accuracy.
  • Insufficient and undifferentiated supervision signals. Most models using both explicit and implicit feature interaction methods require a fusion layer to obtain the final prediction (Wang et al., 2017, 2021b; Mao et al., 2023). However, they only use the final prediction to compute the loss, rather than providing appropriate supervision signals for the different methods themselves. This weakens the effectiveness of the supervision signals. Additionally, some studies (Wang et al., 2023b; Liu et al., 2024), such as the CL4CTR (Wang et al., 2023b) in Figure 1, attempt to introduce auxiliary loss to provide extra supervision signals. However, this often introduces additional computational costs and loss trade-off hyperparameters, increasing the difficulty of hyperparameter tuning. Therefore, a simple, general, yet effective method for computing supervision signals is crucial.
To address the aforementioned limitations, this paper proposes the next generation Deep Cross Network (DCNv3) and the Shallow & Deep Cross Network (SDCNv3), which integrates both low-order and high-order feature interactions while ensuring model interpretability by avoiding the use of DNN for implicit high-order feature interaction modeling. Specifically, we introduce a new exponentially growing Deep Crossing method to explicitly model high-order feature interactions, and we design a Self-Mask operation to filter noise and reduce the number of parameters in the cross network by half. In the fusion layer, we propose a simple yet effective multi-loss trade-off strategy and calculation method, called Tri-BCE, to provide suitable supervision signals for different sub-networks.
The core contributions of this paper are summarized as follows:
  • To the best of our knowledge, this is the first work to achieve surprising performance using only explicit feature interaction modeling without integrating DNN, which may contrast with the popular paradigms in the past CTR prediction literature.
  • We introduce a novel feature interaction modeling method, called Deep Crossing, which grows exponentially with the number of layers to achieve a genuine deep cross network (DCNv3). This method explicitly captures feature interaction information while using a Self-Mask operation to reduce the number of parameters by half and filter noise.
  • We propose a novel CTR model, called SDCNv3, that explicitly captures both low-order and high-order feature interactions. Additionally, we introduce a simple yet effective multi-loss trade-off and calculation method, called Tri-BCE, to ensure that different sub-networks receive appropriate supervision signals.
  • Comprehensive experiments on six datasets demonstrate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. Based on our experimental results, our models achieve 1st rankings on multiple CTR prediction benchmarks.
2. Related Work and Background
2.1. CTR Prediction
Effectively capturing feature interactions has always been one of the key methods for improving CTR prediction, thus receiving extensive research attention (Tian et al., 2023; Wang et al., 2023a; Zhu et al., 2023b). Traditional methods include LR (Richardson et al., 2007), which captures first-order feature interactions, and FM (Rendle, 2010) and its derivatives (Sun et al., 2021; Xiao et al., 2017; Pan et al., 2018), which capture second-order feature interactions. With the rise of deep learning, several models attempt to use DNN to capture higher-order feature interactions (e.g., PNN (Qu et al., 2016), Wide & Deep (Cheng et al., 2016), DeepFM (Guo et al., 2017), DCNv1 (Wang et al., 2017), DCNv2 (Wang et al., 2021b), and DIN (Zhou et al., 2018)), achieving better performance. Among these, the DCN series models are widely recognized for their effective trade-off between efficiency and performance, gaining significant attention from both academia and industry (Wang et al., 2017, 2021b; Chen et al., 2021; Wang et al., 2023a; Yu et al., 2021; Lian et al., 2018). Most subsequent deep CTR models follow the paradigm established by DCN, integrating explicit and implicit feature interactions.
Explicit feature interactions are often modeled directly through hierarchical structures, such as the Cross Layer in the DCN (Wang et al., 2017), the Graph Layer in FiGNN (Li et al., 2019), and the Interacting Layer in AutoInt (Song et al., 2019). These methods ensure partial interpretability while allowing the capture of higher-order feature interactions. On the other hand, implicit feature interactions use DNN to automatically learn complex, non-manually defined patterns and interactions among features, further enhancing model performance (Wang et al., 2021a; Cheng et al., 2020; Wang et al., 2021b).
However, as the performance of explicit feature interactions is generally weaker than that of implicit feature interactions (Mao et al., 2023), several models attempt to abandon standalone explicit interaction methods and instead integrate multiplicative operations into DNN. MaskNet (Wang et al., 2021a) introduces multiplicative operations block by block, while GateNet (Huang et al., 2020), PEPNet (Chang et al., 2023), and FINAL (Zhu et al., 2023b) introduce them layer by layer to achieve higher performance. However, solely pursuing implicit modeling leads to a lack of interpretability. Meanwhile, most models lack the ability to filter noise and obtain appropriate supervisory signals. This paper aims to address these limitations through our proposed methods.
2.2. Problem Definition
2.2.1. DEFINITION 1: CTR Prediction.
It is typically considered a binary classification task that utilizes user profiles (Zhu et al., 2021; Song et al., 2019), item attributes, and context as features to predict the probability of a user clicking on an item. The composition of these three types of features is as follows:
  • User profiles ($x_{U}$): age, gender, occupation, etc.
  • Item attributes ($x_{I}$): brand, price, category, etc.
  • Context ($x_{C}$): timestamp, device, position, etc.
Further, we can define a CTR input sample in the tuple data format: $X=\{x_{U},x_{I},x_{C}\}$. Variable $y\in\{0,1\}$ is an true label for user click behavior:
(1) $y=\begin{cases}1,&\text{user}\text{ has clicked }\text{item},\\ 0,&\text{otherwise. }\end{cases}$
A positive sample when $y=1$ and a negative sample when $y=0$. A CTR prediction model aims to predict $y$ and rank items based on the predicted probabilities $\hat{y}$.
2.2.2. DEFINITION 2: Feature Interaction.
Implicit feature interaction aims to automatically learn complex non-manually defined data patterns and high-order feature interactions using DNN. It is characterized by high efficiency, strong performance, and poor interpretability. Additionally, (Rendle et al., 2020) points out the inefficiency of DNN in learning multiplicative operations.
Explicit feature interaction aims to model the combinations and relationships between input features directly through predefined functions, thereby improving model interpretability. A popular explicit feature interaction method is $X_{2}=X_{1}\odot X_{1}$ (Wang et al., 2021a, 2017; He and Chua, 2017; Wang et al., 2021b; Chen et al., 2021). This method uses the Hadamard Product to interact with two first-order features and generate a second-order feature.
3. Proposed Architecture
Figure 2. Architecture of the SDCNv3.
3.1. Embedding & Reshape Layer
The input $X$ of the CTR prediction task, which is multi-field categorical data and is represented using one-hot encoding. Most CTR prediction models (Song et al., 2019; Li et al., 2023; Sang et al., 2024) utilize an embedding layer to transform them into low-dimensional dense vectors: $\mathbf{e}_{i}=\textit{E}_{i}x_{i}$, where $\textit{E}_{i}\in\mathbb{R}^{d\times s_{i}}$ and $s_{i}$ separately indicate the embedding matrix and the vocabulary size for the $i$-th field, $d$ represents the embedding dimension. In our model, to enable the Cross & Masked Vector to share a weight matrix and separately process two different information streams, we reshape the embeddings using the chunk operation to divide them into two different views: $\mathbf{e}_{i,a},\mathbf{e}_{i,b}=\texttt{chunk}(\mathbf{e}_{i})$. Further, we can get the resulting embedding after reshaping: $\mathbf{x_{1}}=\left[\mathbf{e}_{1,a},\cdots,\mathbf{e}_{f,a},\mathbf{e}_{1,b},\cdots,\mathbf{e}_{f,b}\right]\in\mathbb{R}^{D}$, where $f$ denotes the number of fields, $D=\sum_{i=1}^{f}d$, and $\mathbf{x_{1}}$ denotes the first-order features.
3.2. Shallow & Deep Cross Network v3
The architecture of SDCNv3, shown in Figure 2, integrates different Crossing methods and various layers to simultaneously capture both high-order and low-order explicit feature interactions. As the core structure of SDCNv3, the shallow explicit modeling of SCNv3 recursive formula is as follows:
(2) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{l}+\mathbf{b}_{l},$
$\mathbf{x}_{l+1}$ $=\mathbf{x}_{1}\odot\left[\mathbf{c}_{l}\ ||\ \texttt{Mask}(\mathbf{c}_{l})\right]+\mathbf{x}_{l},$
where $\mathbf{c}_{l}\in\mathbb{R}^{\frac{D}{2}}$ represents the Cross Vector at $l\text{-}th$ layer, Mask denotes the Self-Mask operation, $\mathbf{W}_{l}\in\mathbb{R}^{\frac{D}{2}\times D}$ and $\mathbf{b}_{l}\in\mathbb{R}^{\frac{D}{2}}$ are the learnable weight matrix and bias vector, respectively, and $\mathbf{x}_{l}\in\mathbb{R}^{D}$ represents the $l\text{-}th$ order feature interaction.
Figure 3. Visualization of different crossing methods. $\times$ denotes matrix multiplication. The left side shows the workflow of Cross Network v2 (Wang et al., 2021b), while the right side shows the workflow of Shallow & Deep Cross Network v3. To simplify the representation and facilitate understanding of their differences, we use the same math notation.
The deep explicit modeling of DCNv3 recursive formula is as follows:
(3) $\mathbf{c}_{l}$ $=\mathbf{W}_{l}\mathbf{x}_{2^{l-1}}+\mathbf{b}_{l},$
$\mathbf{x}_{{2^{l}}}$ $=\mathbf{x}_{2^{l-1}}\odot\left[\mathbf{c}_{l}\ ||\ \texttt{Mask}(\mathbf{c}_{l})\right]+\mathbf{x}_{2^{l-1}},$
where $\mathbf{x}_{2^{l}}\in\mathbb{R}^{D}$ represents the $2^{l}\text{-}th$ order feature interaction. From Figure 3, we can see that the original Cross Network v2 is essentially a shallow crossing method, achieving linear growth in the order of feature interactions through the stacking of layers. Experiments show that the optimal number of cross layers in DCNv2 is 2 or 3 (Wang et al., 2021b), with most of its performance coming from the DNN responsible for implicit interactions. In contrast, DCNv3 modifies $\mathbf{x}_{1}$ to $\mathbf{x}_{2^{l-1}}$ to achieve exponential growth in the order of feature interactions.
Figure 3 visualizes the computation process of different crossing methods. The Cross Vector $\mathbf{c}_{l}$ aims to model interactions between features at the bit-level, while the weight matrix $\mathbf{W}_{l}$ aims to compute the inherent importance of different feature fields. However, as pointed out by some works (Xiao et al., 2017; Liu et al., 2020; Wang et al., 2023a), not all feature interactions are beneficial for the final prediction in CTR tasks. Therefore, we introduce the Self-Mask operation to filter out noisy information from feature interactions in another view while maintaining the integrity of the interaction information of the original view. To avoid additional computational costs, we use LayerNorm (Ba et al., 2016) with element-wise affine to regularize it, ensuring a mask rate of around 0.5. The specific formalization of the Self-Mask operation is as follows:
(4) $\texttt{Mask}(\mathbf{c}_{l})=\mathbf{c}_{l}\odot\texttt{max}(0,\texttt{LayerNorm}(\mathbf{c}_{l})).$
Other mask mechanisms can also be used here, such as random Mask based on the Bernoulli distribution, learnable Mask based on Top-K selection, etc. To ensure our proposed model is simple and effective, we use LayerNorm (Ba et al., 2016) to perform a straightforward and efficient normal distribution transformation on $\mathbf{c}_{l}$, ensuring that its output contains approximately 50% zero values to filter out noise, and it reduces the number of parameters by half compared to Cross Network v2.
3.3. Fusion Layer
Most previous CTR models attempt to capture both implicit and explicit feature interactions, which essentially means capturing low-order and high-order feature interactions. Our SDCNv3 achieves this only through explicit modeling. On the other hand, existing works (Wang et al., 2021b, 2023a; Zhu et al., 2023b) often integrate sub-networks using either parallel or stacked structures. Considering the high parallelism of the former, we use a parallel structure to fuse information and compute the loss:
(5) $\hat{y}_{\scriptscriptstyle D}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle D}\mathbf{x}_{{2^{L}}}+\mathbf{b}_{\scriptscriptstyle D}),$
$\hat{y}_{\scriptscriptstyle S}$ $=\sigma(\mathbf{W}_{\scriptscriptstyle S}\mathbf{x}_{L+1}+\mathbf{b}_{\scriptscriptstyle S}),$
$\hat{y}$ $=\texttt{Mean}(\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}),$
where $\mathbf{W}_{\scriptscriptstyle D}$ and $\mathbf{W}_{\scriptscriptstyle S}\in\mathbb{R}^{1\times D}$ represent learnable weights, $\mathbf{b}_{\scriptscriptstyle D}$ and $\mathbf{b}_{\scriptscriptstyle S}$ are biases, Mean denotes the mean operation, $\hat{y}_{\scriptscriptstyle D},\hat{y}_{\scriptscriptstyle S}$ represent the prediction results of DCNv3 and SCNv3, respectively, and $L$ denotes the last number of layers.
Figure 4. The workflow for the Tri-BCE loss.
Tri-BCE loss calculation and trade-off method are shown in Figure 4. We use the widely adopted binary cross-entropy loss (Wang et al., 2023a; Zhu et al., 2023b; Mao et al., 2023) (i.e., Logloss) as both the primary and auxiliary loss for the SDCNv3:
(6) $\mathcal{L}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{i}\right)\right),$
$\mathcal{L}_{\scriptscriptstyle D}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle D,i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle D,i}\right)\right),$
$\mathcal{L}_{\scriptscriptstyle S}$ $=-\frac{1}{N}\sum_{i=1}^{N}\left(y_{i}\log\left(\hat{y}_{\scriptscriptstyle S,i}\right)+\left(1-y_{i}\right)\log\left(1-\hat{y}_{\scriptscriptstyle S,i}\right)\right),$
where $y$ denotes the true labels, $N$ denotes the batch size, $\mathcal{L}_{\scriptscriptstyle D}$ and $\mathcal{L}_{\scriptscriptstyle S}$ represent the individual losses for the prediction results of DCNv3 and SCNv3, respectively, and $\mathcal{L}$ represents the primary loss. To provide each sub-network with suitable supervision signals, we assign them adaptive weights, $\mathbf{w}_{\scriptscriptstyle D}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle D}-\mathcal{L})$ and $\mathbf{w}_{\scriptscriptstyle S}=\texttt{max}(0,\mathcal{L}_{\scriptscriptstyle S}-\mathcal{L})$, and jointly train them to achieve Tri-BCE loss:
(7) $\mathcal{L}_{\text{Tri}}$ $=\mathcal{L}+\mathbf{w}_{\scriptscriptstyle D}\cdot\mathcal{L}_{\scriptscriptstyle D}+\mathbf{w}_{\scriptscriptstyle S}\cdot\mathcal{L}_{\scriptscriptstyle S},$
As demonstrated by (Li et al., 2024), providing a single supervision signal to sub-networks is often suboptimal. Our proposed Tri-BCE loss helps sub-networks learn better parameters by providing adaptive weights that change throughout the learning process. Theoretically, we can derive the gradients obtained by $\hat{y}_{\scriptscriptstyle D}$:
(8) $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle D})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log\hat{y}^{+}+\mathbf{w}_{\scriptscriptstyle D}\log\hat{y}^{+}_{\scriptscriptstyle D}\right)}{\partial\hat{y}^{+}_{\scriptscriptstyle D}}=-\frac{1}{N}\left(\frac{1}{2\hat{y}^{+}}+\frac{\mathbf{w}_{\scriptscriptstyle D}}{\hat{y}^{+}_{\scriptscriptstyle D}}\right),$
$\nabla_{(\hat{y}^{-}_{\scriptscriptstyle D})}\mathcal{L}_{\text{Tri}}$ $=-\frac{1}{N}\cdot\frac{\partial\left(\log(1-\hat{y}^{-})+\mathbf{w}_{\scriptscriptstyle D}\log(1-\hat{y}^{-}_{\scriptscriptstyle D})\right)}{\partial\hat{y}^{-}_{\scriptscriptstyle D}}$
$=\frac{1}{N}\left(\frac{1}{2(1-\hat{y}^{-})}+\frac{\mathbf{w}_{\scriptscriptstyle D}}{1-\hat{y}^{-}_{\scriptscriptstyle D}}\right),$
where $\nabla_{(\hat{y}^{+}_{\scriptscriptstyle D})}$ and $\nabla_{(\hat{y}^{-}_{\scriptscriptstyle D})}$ represent the gradients received by $\hat{y}_{\scriptscriptstyle D}$ for positive and negative samples, respectively. Similarly, the gradient signals received by $\hat{y}_{\scriptscriptstyle S}$ are consistent with those of $\hat{y}_{\scriptscriptstyle D}$, so we do not elaborate further. It can be observed that $\hat{y}_{\scriptscriptstyle D}$ and $\hat{y}_{\scriptscriptstyle S}$ both have the same gradient terms $\frac{1}{2\hat{y}^{+}}$ and $\frac{1}{2(1-\hat{y}^{-})}$, indicating that training both sub-networks with a single loss results in identical supervision signals, which is detrimental to the model’s learning. However, our Tri-BCE loss additionally provides dynamically adjusted gradient terms based on $\mathbf{w}_{\scriptscriptstyle D}$ and $\mathbf{w}_{\scriptscriptstyle S}$, ensuring that the sub-networks are directly influenced by the true labels $y$ and adaptively adjust their weights according to the difference between the primary and auxiliary losses. Therefore, Tri-BCE loss provides the sub-networks with more suitable supervision signals.
3.4. Complexity Analysis
To further compare the time complexity of the DCN series models, we discuss and analyze the time complexity of different models. Let $W_{\Psi}$ denote the predefined number of parameters in the DNN, and $s$ denote the feature vocabulary size. The definitions of the other variables can be found in the previous sections. For clarity, we further provide a comparison of the magnitudes of different variables in Table 1. We can derive:
Table 1. Comparison of Analytical Time Complexity
$s\gg|W_{\Psi}|>D>f\approx d>L$
Model Embedding Implicit interaction Explicit interaction Loss Function
DCNv1 (Wang et al., 2017) O($dfs$) O($|W_{\Psi}|$) O($2DL$) O($N$)
DCNv2 (Wang et al., 2021b) O($dfs$) O($|W_{\Psi}|$) O($D^{2}L$) O($N$)
EDCN (Chen et al., 2021) O($dfs$) O($D^{2}L$) O($D^{2}L$) O($N$)
GDCN (Wang et al., 2023a) O($dfs$) O($|W_{\Psi}|$) O($2D^{2}L$) O($N$)
DCNv3 O($dfs$) - O($\frac{D^{2}L}{2}$) O($N$)
[…]
  • All models have the same time complexity for embedding. Therefore, we only visualize the non-embedding parameters in the experiment section.
  • Except for our proposed DCNv3 and SDCNv3, all other models include implicit interaction to enhance predictive performance, which incurs additional computational costs.
  • In terms of explicit interaction, DCNv3 only has a higher time complexity than DCNv1, and the time complexity of GDCN is four times that of DCNv3.
  • Since our SDCNv3 uses the Tri-BCE loss, the time complexity of loss computation for SDCNv3 is three times that of other models. However, this does not affect the model’s inference speed.
4. Experiments
In this section, we conduct comprehensive experiments on six CTR prediction datasets to validate the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3, and address the following research questions (RQs):
  • RQ1 Do DCNv3 and SDCNv3 outperform other CTR models in terms of performance? Do they perform well on large-scale and highly sparse datasets?
  • RQ2 Are DCNv3 and SDCNv3 more efficient compared to other CTR models?
  • RQ3 Do SDCNv3 possess interpretability and the ability to filter noise?
  • RQ4 How do different configurations affect the models?
4.1. Experiment Setup
4.1.1. Datasets.
We evaluate DCNv3 and SDCNv3 on six CTR prediction datasets: Avazu [^1] https://www.kaggle.com/c/avazu-ctr-prediction (Zhu et al., 2021), Criteo [^2] https://www.kaggle.com/c/criteo-display-ad-challenge (Zhu et al., 2021), ML-1M [^3] https://grouplens.org/datasets/movielens (Song et al., 2019), KDD12444https://www.kaggle.com/c/kddcup2012-track2 (Song et al., 2019), iPinYou [^5] https://contest.ipinyou.com/ (Qu et al., 2018), and KKBox [^6] https://www.kkbox.com/intl (Zhu et al., 2022a). Table 2 provides detailed information about these datasets. A more detailed description of these datasets can be found in the given references and links.
Table 2. Dataset statistics
Dataset #Instances #Fields #Features
Avazu 40,428,967 24 3,750,999
Criteo 45,840,617 39 910,747
ML-1M 739,012 7 9,751
KDD12 141,371,038 13 4,668,096
iPinYou 19,495,974 16 665,765
KKBox 7,377,418 13 91,756
4.1.2. Data Preprocessing.
We follow the approach outlined in (Zhu et al., 2021). For the Avazu dataset, we transform the timestamp field it contains into three new feature fields: hour, weekday, and weekend. For the Criteo and KDD12 dataset, we discretize the numerical feature fields by rounding down each numeric value $x$ to $\lfloor\log^{2}(x)\rfloor$ if $x>2$, and $x=1$ otherwise. We set a threshold to replace infrequent categorical features with a default ”OOV” token. We set the threshold to 10 for Criteo, KKBox, and KDD12, 2 for Avazu and iPinYou, and 1 for the small dataset ML-1M. More specific data processing procedures and results can be found in our open-source run logs [^7] https://anonymous.4open.science/r/DCNv3-E352/checkpoints/ and configuration files, which we do not elaborate on here.
4.1.3. Evaluation Metrics.
To compare the performance, we utilize two commonly used metrics in CTR models: Logloss, AUC (Song et al., 2019; Wang et al., 2023a; Zhu et al., 2022b). AUC stands for Area Under the ROC Curve, which measures the probability that a positive instance will be ranked higher than a randomly chosen negative one. Logloss is the result of the calculation of $\mathcal{L}$ in Equation 6. A lower Logloss suggests a better capacity for fitting the data. It is worth noting that even a slight improvement (e.g., 0.1%) in Logloss and AUC is meaningful in the context of CTR prediction tasks (Chen et al., 2021; Wang et al., 2023b; Zhu et al., 2021).
4.1.4. Baselines.
We compared DCNv3 and SDCNv3 with some state-of-the-art (SOTA) models (* denotes Integrating the original model with DNN networks): (1) Since DCNv3 is a standalone network that performs explicit feature interactions, we compare it with several models that also perform explicit feature interactions on two large-scale datasets. For example, LR (2007) (Richardson et al., 2007) implements first-order feature interactions; FM and its derivative models FM (2010) (Rendle, 2010), FwFM (2018) (Pan et al., 2018), AFM (2017) (Xiao et al., 2017), FmFM (2021) (Sun et al., 2021) implement second-order feature interactions; and CrossNetv1 (2017) (Wang et al., 2017), CrossNetv2 (2021) (Wang et al., 2021b), CIN (2018) (Lian et al., 2018), AutoInt (2019) (Song et al., 2019), AFN (2020) (Cheng et al., 2020), FiGNN (2019) (Li et al., 2019) implement higher-order feature interactions. (2) To verify the superiority of DCNv3 and SDCNv3 over models that include implicit feature interactions, we further select several high-performance representative baselines, such as PNN (2016) (Qu et al., 2016), Wide & Deep (2016) (Cheng et al., 2016), DeepFM (2017) (Guo et al., 2017), DCNv1 (2017) (Wang et al., 2017), xDeepFM (2018) (Lian et al., 2018), AutoInt* (2019) (Song et al., 2019), AFN* (2020) (Cheng et al., 2020), DCNv2 (2021) (Wang et al., 2021b), EDCN (2021) (Chen et al., 2021), MaskNet (2021) (Wang et al., 2021a), CL4CTR (2023) (Wang et al., 2023b), EulerNet (2023) (Tian et al., 2023), FinalMLP (2023) (Mao et al., 2023), and FINAL (2023) (Zhu et al., 2023b).
4.1.5. Implementation Details.
We implement all models using PyTorch (Paszke et al., 2019) and refer to existing works (Zhu et al., 2021; Huawei, 2021). We employ the Adam optimizer (Kingma and Ba, 2014) to optimize all models, with a default learning rate set to 0.001. For the sake of fair comparison, we set the embedding dimension to 128 for KKBox and 16 for the other datasets (Zhu et al., 2021, 2022a). The batch size is set to 4,096 on the Criteo, ML-1M, and iPinYou datasets and 10,000 on the other datasets. To prevent overfitting, we employ early stopping with a patience value of 2. The hyperparameters of the baseline model are configured and fine-tuned based on the optimal values provided in (Huawei, 2021; Zhu et al., 2021) and their original paper. Further details on model hyperparameters and dataset configurations are available in our straightforward and accessible running logs [^8] https://anonymous.4open.science/r/DCNv3-E352/checkpoints/ and are not reiterated here.
4.2. Overall Performance (RQ1)
Table 3. Performance comparison of different explicit CTR models. Abs.Imp represents the absolute performance improvement of DCNv3 over the strongest baseline.
Model Class Models Avazu Criteo
[…]
First-order LR (Richardson et al., 2007) 0.3815 77.74 0.4567 79.33
Second-order FM (Rendle, 2010) 0.3753 78.87 0.4431 80.86
FwFM (Pan et al., 2018) 0.3744 79.06 0.4407 81.12
AFM (Xiao et al., 2017) 0.3792 78.25 0.4455 80.60
FmFM (Sun et al., 2021) 0.3763 78.57 0.4399 81.22
Higher-order CrossNetv1 (Wang et al., 2017) 0.3779 78.39 0.4456 80.59
CrossNetv2 (Wang et al., 2021b) 0.3748 78.82 0.4392 81.27
CIN (Lian et al., 2018) 0.3742 78.94 0.4393 81.27
AutoInt (Song et al., 2019) 0.3745 78.91 0.4399 81.19
AFN (Cheng et al., 2020) 0.3740 79.06 0.4402 81.15
FiGNN (Li et al., 2019) 0.3738 79.11 0.4382 81.38
DCNv3 0.3698 79.65 0.4364 81.55
Abs.Imp -0.004 +0.54 -0.0018 +0.17
4.2.1. Comparison with models using only explicit feature interactions.
Since DCNv3 models feature interactions explicitly, we select 11 representative models for comparison, categorized into First-order, Second-order, and Higher-order classes. We bold the best performance, while underlined scores are the second best. The experimental results are shown in Table 3, and we can draw the following conclusions:
  • By comparing Table 3 and Table 4, we find that most models using only explicit feature interactions often perform worse than those integrating implicit feature interactions, and even worse than a simple DNN. This undoubtedly undermines the necessity of explicit feature interactions.
  • Overall, capturing higher-order feature interactions often enhances model performance. For example, FM outperforms LR on the two large-scale datasets Avazu and Criteo, and CrossNetv2 outperforms all first-order and second-order feature interaction models except for FwFM on Avazu. This demonstrates the effectiveness of higher-order feature interactions in improving model performance.
  • More complex model structures do not necessarily lead to performance improvements. AFM introduces a more complex attention mechanism compared to FM, yet it does not achieve better performance, as also reported in (Zhu et al., 2021). However, CrossNetv2 extends the size of the weight matrix compared to CrossNetv1, resulting in a certain degree of performance enhancement. Therefore, we should carefully design the model architecture.
  • FiGNN achieves the best baseline performance among the explicit feature interaction models. However, our DCNv3 still achieves a Logloss decrease of 0.4% and an AUC increase of 0.54% on the Avazu dataset compared to FiGNN, and a Logloss decrease of 0.18% and an AUC increase of 0.17% on the Criteo dataset, both exceeding 0.001 in a statistically significant level. This demonstrates the superiority of DCNv3.
Table 4. Performance comparison of different deep CTR models. ”*”: Integrating the original model with DNN networks. Meanwhile, we conduct a two-tailed T-test to assess the statistical significance between our models and the best baseline ($\star$: $p$ 1e-3). Abs.Imp represents the absolute performance improvement of SDCNv3 over the strongest baseline. Typically, CTR researchers consider an improvement of 0.001 (0.1%) in Logloss and AUC to be statistically significant (Wang et al., 2017; Chen et al., 2021; Wang et al., 2023b; Zhu et al., 2021).
  Avazu Criteo ML-1M KDD12 iPinYou KKBox
[…]
DNN (Covington et al., 2016) 0.3726 79.18 0.4380 81.40 0.3100 90.30 0.1502 80.52 0.005545 78.06 0.4811 85.01
PNN (Qu et al., 2016) 0.3719 79.32 0.4378 81.42 0.3070 90.42 0.1504 80.47 0.005544 78.13 0.4793 85.15
Wide & Deep (Cheng et al., 2016) 0.3725 79.20 0.4376 81.42 0.3056 90.45 0.1504 80.48 0.005542 78.09 0.4852 85.04
DeepFM (Guo et al., 2017) 0.3723 79.21 0.4375 81.43 0.3073 90.51 0.1501 80.60 0.005549 77.94 0.4785 85.31
DCNv1 (Wang et al., 2017) 0.3725 79.21 0.4376 81.44 0.3156 90.38 0.1501 80.59 0.005541 78.13 0.4766 85.31
xDeepFM (Lian et al., 2018) 0.3722 79.24 0.4376 81.43 0.3054 90.47 0.1501 80.62 0.005534 78.25 0.4772 85.35
AutoInt* (Song et al., 2019) 0.3722 79.24 0.4390 81.32 0.3112 90.45 0.1502 80.57 0.005544 78.16 0.4773 85.34
AFN* (Cheng et al., 2020) 0.3727 79.21 0.4384 81.38 0.3048 90.53 0.1499 80.70 0.005539 78.17 0.4842 84.89
DCNv2 (Wang et al., 2021b) 0.3724 79.22 0.4376 81.45 0.3098 90.56 0.1502 80.59 0.005539 78.26 0.4787 85.31
EDCN (Chen et al., 2021) 0.3716 79.35 0.4386 81.36 0.3073 90.48 0.1501 80.62 0.005573 77.93 0.4952 85.27
MaskNet (Wang et al., 2021a) 0.3716 79.36 0.4387 81.34 0.3080 90.34 0.1498 80.79 0.005556 77.85 0.5003 84.79
CL4CTR (Wang et al., 2023b) 0.3724 79.21 0.4383 81.35 0.3074 90.33 0.1502 80.56 0.005543 78.06 0.4972 83.78
EulerNet (Tian et al., 2023) 0.3723 79.22 0.4379 81.47 0.3050 90.44 0.1498 80.78 0.005540 78.30 0.4922 84.27
FinalMLP (Mao et al., 2023) 0.3718 79.35 0.4373 81.45 0.3058 90.52 0.1497 80.78 0.005556 78.02 0.4822 85.10
FINAL (Zhu et al., 2023b) 0.3712 79.41 0.4371 81.49 0.3035 90.53 0.1498 80.74 0.005540 78.13 0.4795 85.08
DCNv3 0.3698 79.65 0.4364 81.55 0.3013 90.59 0.1496 80.90 0.005534 78.43 0.4778 85.40
SDCNv3 0.3695 79.70 0.4358 81.62 0.3001 90.74 0.1494 80.98 0.005535 78.56 0.4765 85.57
Abs.Imp -0.0017 +0.29 -0.0013 +0.13 -0.0034 +0.18 -0.0003 +0.19 -0.000004 +0.26 -0.0001 +0.22
Figure 5. Efficiency comparisons with other models on the Criteo dataset. We only consider non-embedding parameters. We fix the optimal performance hyperparameters for each model and conduct experiments uniformly on one GeForce RTX 4090 GPU.
4.2.2. Comparison with models integrating implicit feature interactions.
To further comprehensively investigate the performance superiority and generalization ability of DCNv3 and SDCNv3 on various CTR datasets (e.g., large-scale sparse datasets), we select 15 representative baseline models and 6 benchmark datasets. We highlight the performance of DCNv3 and SDCNv3 in bold and underline the best baseline performance. Table 4 presents the experimental results, from which we can make the following observations:
  • Overall, SDCNv3 achieves the best performance across all six datasets. SDCNv3 shows an average AUC improvement of 0.22% over the strongest baseline model and an average Logloss decrease of 0.11%, both exceeding a statistically significant performance improvement of 0.1%. This demonstrates the effectiveness of SDCNv3. SDCNv3 ranks 1st in CTR prediction on the PapersWithCode benchmarks [^9] https://paperswithcode.com/task/click-through-rate-prediction for the Criteo, KDD12, and KKBox datasets.
  • The FinalMLP model achieves good performance on the Avazu and Criteo datasets, surpassing most CTR models that combine explicit and implicit feature interactions. This demonstrates the effectiveness of implicit feature interactions. Consequently, most CTR models attempt to integrate DNN into explicit feature interaction models to enhance performance. However, SDCNv3 achieves state-of-the-art performance using only explicit feature interactions, indicating the effectiveness of modeling with explicit feature interactions alone.
  • SDCNv3 achieves performance improvements over DCNv3 across all six datasets, demonstrating the effectiveness of SCNv3 in capturing low-order feature interactions and the Tri-BCE loss. Notably, on the iPinYou dataset, we observe that all models have Logloss values around the 0.0055 level. This is due to the imbalance between positive and negative samples in the dataset (Lin et al., 2024), and other works have reported similar results (Qu et al., 2016, 2018).
  • DCNv3 outperforms all baseline models in terms of AUC, with the only exception being Logloss optimization on the KKBox dataset, which is weaker than DCNv1. This further demonstrates the effectiveness of DCNv3, as it captures high-quality feature interaction information through exponentially growing feature interactions and noise filtering mechanisms.
4.3. In-Depth Study of DCNv3 and SDCNv3
4.3.1. Efficiency Comparison (RQ2)
To verify the efficiency of DCNv3 and SDCNv3, we fix the optimal hyperparameters of the 25 baseline models and compare their parameter count (rounded to two decimal places) and runtime (averaged over five runs). The experimental results are shown in Figure 5. We can derive:
  • Explicit CTR models typically use fewer parameters. For instance, LR, FM, FwFM, and AFM have nearly zero non-embedding parameters, while FmFM, CrossNet, CIN, and AutoInt all require fewer than 1M parameters. Notably, parameter count does not always correlate with time complexity. Although CIN uses only 0.57M parameters, its training time per epoch reaches a maximum of 606 seconds, making it unsuitable for practical production environments. FiGNN faces the same issue.
  • The LR model has the lowest runtime at 150 seconds among all models. CrossNetv1 and CrossNetv2 follow closely, requiring only a negligible increase in time while significantly improving performance. This demonstrates the efficiency of CrossNet and its series models. As a fundamental component of deep CTR models, DNN requires only 156 seconds. Due to the parallel-friendly nature of CTR parallel structures, some carefully designed deep CTR models, such as PNN, DCNv2, FinalMLP, and FINAL, significantly enhance prediction accuracy without a substantial increase in runtime.
  • Our proposed DCNv3 and SDCNv3 are the most parameter-efficient models in the DCN series, requiring only 0.78M and 1.37M parameters respectively to achieve SOTA performance. Meanwhile, regarding runtime, DCNv3 consistently outperforms strong baseline models such as FinalMLP, FINAL, DCNv2, and DCNv1. This demonstrates the time-efficient of DCNv3. Although SDCNv3 requires an additional 20 seconds compared to DCNv3 due to the employment of the Tri-BCE loss, it still remains comparable to DCNv2. Notably, the extra computational cost for the loss is only incurred during training and does not affect inference speed in practical applications. This further demonstrates the efficiency of DCNv3 and SDCNv3.
4.3.2. The Interpretability and noise filtering capability of SDCNv3 (RQ3)
Interpretability is a crucial aspect of CTR prediction tasks (Li et al., 2020; Wang et al., 2023a, 2021b; Song et al., 2019), as it helps researchers understand the reasons behind predictions and increases confidence in the results. In this section, we investigate the dynamic Cross & Masked Vector and the static $\mathbf{W}_{l}$ to understand the model’s prediction process. The experimental and visualization results are shown in Figure 6 (Value denotes the Frobenius norm for each feature field), we can derive the following observations:
(a) $\mathbf{c}_{l}$ of SCNv3
(b) $\texttt{Mask}(\mathbf{c}_{l})$ of SCNv3
(c) $\mathbf{c}_{l}$ of DCNv3
(d) $\texttt{Mask}(\mathbf{c}_{l})$ of DCNv3
(e) $\mathbf{W}_{3}$ of SCNv3
(f) $\mathbf{W}_{3}$ of DCNv3
Figure 6. Visualization of field-wise value and sparsity for $\mathbf{c}_{l}$, $\texttt{Mask}(\mathbf{c}_{l})$, and $\mathbf{W}_{3}$ of SDCNv3 on a batch of ML-1M.
[…]
  • $\mathbf{c}_{l}$ and $\texttt{Mask}(\mathbf{c}_{l})$ exhibit complementary properties. When a feature field in $\mathbf{c}_{l}$ becomes important, its corresponding sparsity in $\texttt{Mask}(\mathbf{c}_{l})$ decreases, and vice versa (e.g., UserID, Occupation, Age fields). This demonstrates the effectiveness of our introduced Self-Mask operation, which further filters out noise by more aggressively assigning zero values to the representation elements of certain feature fields.
  • From Figure 6 (e, f), we observe that SCNv3 and DCNv3 capture different feature interaction information at the same layer. In SCNv3, $\mathbf{W}_{3}$ is used to compute the importance of 3-order features to generate 4-order feature interactions. In contrast, in DCNv3, $\mathbf{W}_{3}$ is used to compute the importance of $2^{2}$-order features to generate $2^{3}$-order feature interactions. Consequently, DCNv3 shows reduced importance for UserID $\times$ Genres compared to SCNv3. This further proves the validity of SDCNv3.
  • Overall, we observe that the importance of higher-order feature interactions is lower than that of lower-order feature interactions, which is similarly reported in some works (Liu et al., 2020; Wang et al., 2023a; Zhu et al., 2023a). For example, in Figure 6, (f) has fewer bright red blocks compared to (e), and the blue blocks in (a) gradually darken as the number of layers increases, the situation is similar in (b).
4.3.3. Ablation Study (RQ4)
To investigate the impact of each component of SDCNv3 on its performance, we conduct experiments on several variants of SDCNv3:
  • DCNv3: SDCNv3 without the SCNv3.
  • SCNv3: SDCNv3 without the DCNv3.
  • w/o TB: SDCNv3 with BCE instead of the Tri-BCE.
  • w/o LN: Self-Mask without the LayerNorm.
The results of the ablation experiments are shown in Table 5. It is observed that both DCNv3 and SCNv3 exhibit some performance loss compared to SDCNv3, which demonstrates the necessity of capturing both high-order and low-order feature interactions. Meanwhile, the variant w/o TB also leads to a certain degree of performance decline, particularly noticeable on KKBox. LayerNorm aims to ensure that the Self-Mask maintains a masking rate of around 0.5, so its removal also results in some performance loss. This demonstrates the necessity and effectiveness of each component within SDCNv3.
Table 5. Ablation study of SDCNv3.
  Model Avazu Criteo ML-1M KKBox
[…]
SDCNv3 0.3695 79.70 0.4358 81.62 0.3001 90.74 0.4765 85.57
DCNv3 0.3698 79.65 0.4364 81.55 0.3013 90.59 0.4778 85.40
SCNv3 0.3707 79.49 0.4370 81.50 0.3028 90.54 0.4827 85.24
w/o TB 0.3698 79.64 0.4360 81.57 0.3018 90.62 0.5125 82.43
w/o LN 0.3699 79.63 0.4361 81.58 0.3026 90.66 0.4771 85.52
(a) Criteo
(b) KDD12
Figure 7. Performance comparison for different network depths of DCNv3.
4.3.4. Influence of Network Depths (RQ4)
To further investigate the Influence of different neural network depths on the performance of DCNv3, we conduct experiments on two large-scale CTR datasets, Criteo and KDD12. Figure 7 shows the AUC and Logloss performance of DCNv3 on the test sets.
From Figure 7, we observe that on the Criteo dataset, the model achieves optimal performance at a depth of 4 layers, indicating that DCNv3 captures up to $2^{4}$-order feature interactions. On the KDD12 dataset, DCNv3 achieves optimal performance at a depth of 6 layers, meaning it captures $2^{6}$-order feature interactions. In contrast, achieving the same order of feature interactions in the linearly growing CrossNetv2 requires $12^{4}-$ 1 and $12^{6}-$ 1 layers, respectively. Considering the huge computational resources required, this is impractical, whereas DCNv3 easily accomplishes this with its exponentially growing feature interaction mechanism. This further demonstrates the effectiveness of DCNv3.
5. Conclusion
This paper introduced the next generation deep cross networks, DCNv3 and SDCNv3. The former explicitly captured feature interaction through an exponentially growing modeling method, called Deep Crossing, and further filtered noise signals via the Self-Mask operation, reducing the parameter count by half. The latter built on DCNv3 by incorporating the shallow cross network, called SCNv3, to capture both high-order and low-order feature interactions without relying on the less interpretable DNN. Tri-BCE helped the two sub-networks in SDCNv3 obtain more suitable supervision signals for themselves. Comprehensive experiments on six datasets demonstrated the effectiveness, efficiency, and interpretability of DCNv3 and SDCNv3. Additionally, our proposed models achieved 1st rankings in multiple CTR benchmarks using only explicit feature interactions, breaking the convention that traditional CTR models must integrate implicit feature interactions to improve performance.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.